Hello,

For question three I do not know how to start the problem. But for question 4 I wanted to make sure that the answer was correct.

Thank you!

YS 231-01 Spring 2019 Chapter 20 homework 3. Suppose we shoot a beam of unknown charged particles through an electric field of 100 N/C. Fh and the trajectory of the beam, we determined that its acceleration is 1.75x1013 m/s?. What is the c ratio (q/m) of this particle? From the input velocity ae to mass Set up (2 points): Calculation (2 points): 5E13 Answer (1 point): Consider a molecule with displaced charge of 1.6x10-19 C and a separation of 5x10-12 m. If this molecule is placed inside of an electric field with strength 1000 N/C, what would be the maximum torque acting on it? Set up (2 points): 4. q: ).GE-IqC > Calculation (2 points): Answer (1 point): しMAx® 8E-28 N·m
Answer #1

your both answers are correct.

for 1st question

electric force = mass* acceleration

but electric force = q*E

so, q*E=m*a

q/m = a/E = 1.75*E13/100 =1.75*E12 (ans)

and question 2 is also correct